### v1

# Explicit quadratic large sieve inequality

Abstract.
In this article, we obtain an explicit version of Heath-Brown’s large sieve inequality for quadratic characters and discuss its applications to $L$-functions and quadratic fields.
2020 Mathematics Subject Classification:
Primary 11L40, 11N35; Secondary 11L26, 11N36
1. Introduction
1.1. Motivations and applications
In many instances of analytic number theory (e.g. §7, §12, and §17 in [5]), one is confronted with the task of estimating mean squares of character sums
$A(\mathcal{S},N,a)=\sum_{\chi\in\mathcal{S}}\left|\sum_{n\leq N}a_{n}\chi(n)\right|^{2},$
where $\mathcal{S}$ is a designated finite collection of Dirichlet characters, $N\geq 1$, and $a=\{a_{n}\}$ is any sequence of complex numbers. By the Cauchy–Schwartz inequality, one has $A(\mathcal{S},N,a)\leq|\mathcal{S}|N\|a\|^{2}$, so
$\Delta(\mathcal{S},N)=\sup_{\|a\|\neq 0}{A(\mathcal{S},N,a)\over\|a\|^{2}}<\infty.$
A preliminary method to improve $\Delta(\mathcal{S},N)$ makes use of the orthogonality relations of Dirichlet characters. Expanding $A(\mathcal{S},N,a)$ gives
$A(\mathcal{S},N,a)=\sum_{n\leq N}|a_{n}|^{2}\sum_{\chi}|\chi(n)|^{2}+\sum_{\chi}\sum_{n_{1}\neq n_{2}}a_{n_{1}}\overline{a_{n_{2}}}\chi(n_{1})\overline{\chi}(n_{2}).$
The first “diagonal” term is $\leq|\mathcal{S}|\|a\|^{2}$, and the second “off-diagonal” term is usually treated with the orthogonality of characters in $\mathcal{S}$. When $\mathcal{S}$ denotes all characters modulo $q$, this method gives (see [8, Theorem 6.2])
$\Delta(\mathcal{S},N)\leq q+N,$
and when $\mathcal{S}$ denotes all primitive characters with conductor $\leq Q$, there is
(1) $\Delta(\mathcal{S},N)\leq Q^{2}+N-1,$
which is known as the large sieve inequality [1, §27].
In a fundamental paper, Heath-Brown [4, Corollary 2] investigated the case where $\mathcal{S}$ is $S(Q)$, the set of all primitive quadratic characters with conductor $\leq Q$ and proved that for all $\varepsilon>0$, we have
(2) $A(S(Q),N,a)\ll_{\varepsilon}(QN)^{\varepsilon}(Q+N)\sum_{\begin{subarray}{c}n_{1},n_{2}\leq N\\ n_{1}n_{2}=\square\end{subarray}}|a_{n_{1}}a_{n_{2}}|,$
which is referred to as the quadratic large sieve inequality. The estimate (2) is optimal up to the magnitude of the $\ll_{\varepsilon}$-constant.
Compared to the large sieve, the inequality (2) achieved a power saving in $Q$-aspect by relaxing the $\ell^{2}$-norm $\|a\|$. When $\{a_{n}\}$ is supported on squarefree integers, $\sum_{n_{1}n_{2}=\square}|a_{n_{1}}a_{n_{2}}|$ coincides with $\|a\|^{2}$. For general $\{a_{n}\}$, the quantity can be conveniently estimated. For instance, according to [4, p. 267], there is
(3) $\sum_{\begin{subarray}{c}n_{1},n_{2}\leq N\\ n_{1}n_{2}=\square\end{subarray}}|a_{n_{1}}a_{n_{2}}|\ll(N\log N)\max_{n\leq N}|a_{n}|^{2},$
transforming the inequality (2) into a more applicable form.
We cannot replace $\sum_{n_{1}n_{2}=\square}|a_{n_{1}}a_{n_{2}}|$ in (2) with $\|a\|^{2}$ when $\{a_{n}\}$ is an arbitrary sequence. If $a_{n}=\mathbbm{1}_{\square}(n)$, then
$A(S(Q),N,\mathbbm{1}_{\square})$ $=\sum_{\chi\in S(Q)}\Biggl{|}\sum_{\begin{subarray}{c}n^{2}\leq N\\ (n,q_{\chi})=1\end{subarray}}1\Biggr{|}^{2}\geq\sum_{2<p\leq Q}\Biggl{|}\sum_{\begin{subarray}{c}n^{2}\leq N\\ n\neq p\end{subarray}}1\Biggr{|}^{2}\gg{QN\over\log Q},$
but $\|\mathbbm{1}_{\square}\|^{2}\leq\sqrt{N}$, so $A(S(Q),N,a)\ll_{\varepsilon}(QN)^{\varepsilon}(Q+N)\|a\|^{2}$ is generally false.
The quadratic large sieve has produced Lindelöf-on-average-type results for various moments of $L$-functions with quadratic twists. For instance, if $L(s,\chi)$ denotes the Dirichlet $L$-function, Heath-Brown [4, Theorem 2] showed that
(4) $\sum_{\chi\in S(Q)}|L(1/2+it,\chi)|^{4}\ll_{\varepsilon}(Q(|t|+1))^{1+\varepsilon}.$
If $f$ is an $SL_{2}(\mathbb{Z})$-cusp form of weight $\kappa$ that is also an eigenfunction to all Hecke operators with Fourier expansion
$f(z)=\sum_{n\geq 1}\lambda_{f}(n)n^{\kappa-1\over 2}e(nz)$
and $L(s,f\otimes\chi)$ be its $L$-function twisted by $\chi$:
$L(s,f\otimes\chi)=\sum_{n\geq 1}{\lambda_{f}(n)\chi(n)\over n^{s}},$
then an argument similar to the proof of (4) leads to
(5) $\sum_{\chi\in S(Q)}|L(1/2,f\otimes\chi)|^{2}\ll_{\varepsilon}Q^{1+\varepsilon}.$
In 2010, Soundararajan and Young [13] indicated that under GRH, when $\kappa\equiv 0\pmod{4}$ and $S^{*}(Q)$ denotes the collection of $\chi(\cdot)=\left(2\over\cdot\right)\left(\cdot\over m\right)$ with $1\leq m\leq Q$ squarefree odd, there is an asymptotic formula:
(6) $\sum_{\chi\in S^{*}(Q)}L(1/2,f\otimes\chi)^{2}\sim C_{f}Q\log Q.$
The quadratic large sieve inequality (2) has recently gained renewed interest due to the breakthrough of Xiannan Li [7]. Specifically, by improving (2) for special choices of $a=\{a_{n}\}$, Li dropped the requirement of GRH for the asymptotic formula (6), which consequently turns many previously conditional results unconditional.
Apart from moments of $L$-functions with quadratic twists, the quadratic large sieve has been applied to study the frequency of sign changes of Fourier coefficients of half-integral weight modular forms along fundamental discriminants. Specifically, let $k\geq 2$ be an integer and $g(z)$ be a weight $k+\frac{1}{2}$ Hecke eigenform for the group $\Gamma_{0}(4)$ with Fourier expansion
$g(z)=\sum_{n\geq 1}c(n)n^{k-\frac{1}{2}\over 2}e(nz).$
Then Lester and Radziwiłł [6] deduced from (5) that when $N_{g}$ is the number of sign changes of the sequence $\{c(n)\}_{n\in\operatorname{supp}(c)}$, one has
(7) $N_{g}\gg X^{1-\varepsilon}.$
Moreover, if one assumes GRH, this can be improved to $N_{g}\gg X$.
The inequality (2) also has applications to quadratic fields. For instance, let $f(x)\in\mathbb{Z}[x]$ be a fixed separable polynomial of degree $\geq 2$ and $\mathcal{F}(N)$ be the set of reduced fractions $a/b$ such that $1\leq a,b\leq N$. For squarefree number $m\geq 1$, define $R_{f}(m,N)$ to be the number of $r\in\mathcal{F}(N)$ such that $\mathbb{Q}(\sqrt{f(r)})=\mathbb{Q}(\sqrt{m})$. Then Shparlinski [12] applied Heath-Brown’s square sieve [3] and the quadratic large sieve to deduce that
(8) $\sum_{m\leq M}\mu^{2}(m)R_{f}(m,N)\ll_{\varepsilon}N^{4/3+\varepsilon}M^{5/6},$
which improves the trivial bound $O(N^{2})$ in the range $M\leq N^{4/5}$.
For a prime $q\equiv 3\pmod{4}$, denote by $P_{q}(X)$ the number of rational prime $p\leq X$ that splits in $\mathbb{Q}(\sqrt{-q})$. Then Dunn et al. [2] used the quadratic large sieve (2) to show that for all $\varepsilon>0$, the asymptotic
(9) $P_{q}(X)\sim\frac{1}{2}\cdot{X\over\log X}$
is true for almost all $q\sim Q$ provided that $X\geq Q^{\varepsilon}$ and $Q\to+\infty$.
1.2. Statement of results
By the unique factorization of $n$ into $kr^{2}$ with $k$ squarefree and the well-known decomposition of quadratic characters [1, §5], Heath-Brown deduced the estimate (2) by proving the following [4, Theorem 1]:
(10) $\sideset{}{{}^{*}}{\sum}_{m\leq M}\left|\sideset{}{{}^{*}}{\sum}_{n\leq N}a_{n}\left(\frac{n}{m}\right)\right|^{2}\ll_{\varepsilon}(MN)^{\varepsilon}(M+N)\|a\|^{2},$
where $\sum^{*}$ means a sum over squarefree odd integers.
In this paper, we establish an explicit form of this inequality:
Theorem 1.
Let $M,N\geq 1$ and $\{a_{n}\}_{n\sim N}$ be a sequence of complex numbers. Define
$\Sigma(M,N,a)=\sideset{}{{}^{*}}{\sum}_{m\sim M}\left|\sideset{}{{}^{*}}{\sum}_{n\sim N}a_{n}\left(\frac{n}{m}\right)\right|^{2},\quad\mathcal{B}(M,N)=\sup_{\|a\|\neq 0}{\Sigma(M,N,a)\over\|a\|^{2}}.$
Then there exists $C>0$ such that for every $\varepsilon>0$, we have
(11) $\mathcal{B}(M,N)\leq\exp_{4}(C\varepsilon^{-1})\cdot(MN)^{\varepsilon}(M+N).$
This improvement exhibits the explicit $\varepsilon$-dependence in the implied constants, and setting $\varepsilon=C_{1}/\log_{4}X$ for some large $C_{1}>0$ allows us to replace each $X^{\varepsilon}$-term in the estimates (2), (4), (7), (8), and (9) with $\exp(C_{2}\log X/\log_{4}X)$ for some $C_{2}>C_{1}$ when $X\geq\exp_{4}(1)$. To demonstrate this refinement, we prove the following:
Corollary 1 (Explicit quadratic large sieve inequality).
There exists $C>0$ such that for all $Q,N\geq 1$ satisfying $QN\geq\exp_{4}(1)$,
$\sum_{\chi\in S(Q)}\left|\sum_{n\leq N}a_{n}\chi(n)\right|^{2}\leq\exp_{4}(C(QN)/\log_{4}(QN))(Q+N)\sum_{\begin{subarray}{c}n_{1},n_{2}\leq N\\ n_{1}n_{2}=\square\end{subarray}}|a_{n_{1}}a_{n_{2}}|.$
Corollary 2 (Fourth moment of Dirichlet $L$-functions).
There exists $C>0$ such that when $s=\frac{1}{2}+it$, $T=|t|+1$, and $Q\geq\exp_{4}(1)$,
(12) $\sum_{\chi\in S(Q)}|L(s,\chi)|^{4}\leq\exp\{C(QT)/\log_{4}(QT)\}(QT).$
Remark.
When $s=\frac{1}{2}$, Shen and Stucky [11, Corollary 1.2] showed that the right hand side is $\asymp Q(\log Q)^{10}$.
Corollary 3 (Rational primes splitting in $\mathbb{Q}(\sqrt{-q})$).
There exists $C>0$ such that the asymptotic (9) is valid for all but $\ll Q\exp(-C\log Q/\log_{4}Q)$ amount of primes $q\in(Q,2Q]$ that are $\equiv 3\pmod{4}$, provided that $Q\to+\infty$ and $X\geq\exp(3C\log Q/\log_{4}Q)$.
Although our explicit quadratic large sieve (2) can also produce an improvement on (5) analogous to Corollary 2, it is not comparable to the asymptotic obtained by Soundararajan and Young [13] and Xiannan Li [7]. This is because they exploited information concerning the coefficients $\{a_{n}\}$. The former invoked the GRH and estimates on shifted moments of $L$-functions, and the latter utilized the multiplicativity of the coefficients and the approximate functional equations for $L$-functions. In contrast, our Theorem 1 and Corollary 1 make no assumptions on the coefficients $\{a_{n}\}$. This is particularly useful when these coefficients emerge from sieves. For instance, to obtain the estimate (8), Shparlinski applied the quadratic large sieve with coefficients $\{a_{n}\}$ given by Legendre symbols. In Corollary 3, we apply Theorem 1 with $\{a_{n}\}$ supported on primes.
Our Theorem 1 could serve as a baseline for further improvements. As $\exp_{4}(C\varepsilon^{-1})$ still grows rapidly, it would be interesting to see if one can reduce the number of iterations. If the inequality (11) is valid with $\exp(C\varepsilon^{-1})$ instead, then $(MN)^{\varepsilon}$ can be replaced with a fixed positive power of $\log(MN)$, which would be more convenient for various applications.
1.3. Organization of the paper
In Section 2, Section 3, and Section 4, we present outline, lemmas, and the proof for Theorem 1. Then in Section 5, we derive the corollaries.
1.4. Acknowledgements
The work is supported by the LMS Undergraduate Research Bursary URB-2024-41. The author would like to express his sincere gratitude towards Professor Ian Petrow for suggesting the problem and many helpful comments.
1.5. Notations
  • $e(s)$ —– the additive character $e^{2\pi is}$.
  • $n\sim N$ —– the range $N<n\leq 2N$.
  • $\sum_{n}^{*}$ —– the sum over squarefree odd integers $n$.
  • $\|a\|=\left(\sum_{n}|a_{n}|^{2}\right)^{1/2}$ —– the $\ell^{2}$-norm for sequence $a=\{a_{n}\}$.
  • $\exp_{k},\log_{k}$ —– the $k$’th iterated exponential and natural logarithm.
  • $C,C_{1},C_{2},\dots$ —– absolute constants, not necessarily the same in each occurrence.
Unless with a subscript, constants in $O(\cdot)$ and $\ll$ are absolute.
2. Outline for the proof of Theorem 1
As our methods are built on Heath-Brown’s method of recursive estimates, we use the same notations and develop a set of lemmas similar to his. For convenience in bookkeeping, Lemma X in our paper refers to a modification of [4, Lemma X], and Proposition X in this paper refers to a result not present in [4]. Unless stated otherwise, proofs of results mentioned in this section are detailed in Section 4.
We study $\Sigma(M,N,a)$ via the dual sum
$\Sigma_{1}(M,N,a)=\sideset{}{{}^{*}}{\sum}_{m\sim M}\left|\sideset{}{{}^{*}}{\sum}_{n\sim N}a_{n}\left(\frac{m}{n}\right)\right|^{2}.$
Specifically, $\Sigma$ and $\Sigma_{1}$ are related by
Lemma 1.
We have $\mathcal{B}(M,N)\leq 2\mathcal{B}(N,M)$. Moreover, for every sequence $\{a_{n}\}_{n\sim N}$, there exists $\{a_{n}^{\prime}\}_{n\sim N}$ such that $|a_{n}^{\prime}|=|a_{n}|$ and $\Sigma(M,N,a)\leq 2\Sigma_{1}(M,N,a^{\prime})$.
The proof is omitted as the statement is identical to [4, Lemma 1]. From now on, we study $\mathcal{B}(M,N)$ using $\Sigma_{1}$.
2.1. Preparation for Poisson summation
Let
(13) $W(x)=\exp\left(\frac{8}{3}-{2\over 1-(x-\frac{3}{2})^{2}}\right)\chi_{[\frac{1}{2},\frac{5}{2}]}(x).$
Then $W(x)$ is a non-negative smooth function with compact support and $\geq 1$ on $[1,2]$. Although this is different from that used by Heath-Brown [4, p. 241], it still fulfills the purpose and is convenient to derive explicit estimates.
For each integer $m$, let $s(m)$ denote the unique positive squarefree divisor of $m$ that makes $|m/s(m)|$ a square. Then for all $K\leq M/2$, we have
[…]
(14) $\leq\sum_{\begin{subarray}{c}m\text{ odd}\\ s(m)>K\end{subarray}}W\left(\frac{m}{M}\right)\left|\sideset{}{{}^{*}}{\sum}_{n\sim N}a_{n}\left(\frac{m}{n}\right)\right|^{2}:=\Sigma_{1}^{(K)}(M,N,a).$
We expand $\Sigma_{1}^{(K)}$ as follows:
$\Sigma_{1}^{(K)}=\sum_{\begin{subarray}{c}m\text{ odd}\\ s(m)>K\end{subarray}}W\left(\frac{m}{M}\right)\sideset{}{{}^{*}}{\sum}_{n_{1},n_{2}\sim N}a_{n_{1}}\overline{a_{n_{2}}}\left(m\over n_{1}n_{2}\right).$
Let $\Delta=(n_{1},n_{2})$. Then $\chi(m)=\left(m\over n_{1}n_{2}\right)$ is a quadratic Dirichlet character with conductor $q=n_{1}n_{2}/\Delta^{2}$. To utilize this relation, we partition $\Sigma_{1}^{(K)}$ according to the magnitude of $\Delta$. Define
$\Sigma_{2}(M,N,K,\Delta,a)=\sideset{}{{}^{*}}{\sum}_{\begin{subarray}{c}n_{1},n_{2}\sim N\\ (n_{1},n_{2})=\Delta\end{subarray}}a_{n_{1}}\overline{a_{n_{2}}}\sum_{\begin{subarray}{c}m\text{ odd}\\ s(m)>K\end{subarray}}W\left(\frac{m}{M}\right)\left(m\over n_{1}n_{2}\right),$
so we have
$\Sigma_{1}^{(K)}(M,N,a)=\sum_{\Delta\leq 2N}\Sigma_{2}(M,N,K,\Delta,a).$
By making trivial estimates, we can restrict our focus to $\Sigma_{2}$ with small $\Delta$’s:
Lemma 2.
Define
$\mathcal{B}(M,N,K)=\sup_{\|a\|\neq 0}{\Sigma_{1}^{(K)}(M,N,a)\over\|a\|^{2}},\quad\mathcal{C}(M,N,K,\Delta)=\sup_{\|a\|\neq 0}{\Sigma_{2}(M,N,K,\Delta,a)\over\|a\|^{2}}.$
Then there exists absolute $C>0$ such that for all $\varepsilon>0$, $K\leq M/2$, and $\Delta_{0}\leq N$, there is some $N_{1}\leq N/\Delta_{0}$ such that
$\mathcal{B}(M,N,K)\leq\exp_{2}(C\varepsilon^{-1})\cdot N^{\varepsilon}\mathcal{B}(M,N_{1},K)+\sum_{\Delta\leq\Delta_{0}}\mathcal{C}(M,N,K,\Delta).$
To handle $\mathcal{C}(M,N,K,\Delta)$ with $\Delta\leq\Delta_{0}$, we express $\Sigma_{2}$ as a difference:
$\Sigma_{2}(M,N,K,\Delta,a)=\Sigma_{3}(M,N,K,\Delta,a)-\Sigma_{4}(M,N,K,\Delta,a),$
in which
$\Sigma_{3}(M,N,K,\Delta,a)=\sideset{}{{}^{*}}{\sum}_{\begin{subarray}{c}n_{1},n_{2}\sim N\\ (n_{1},n_{2})=\Delta\end{subarray}}a_{n_{1}}\overline{a_{n_{2}}}\sum_{m\text{ odd}}W\left(\frac{m}{M}\right)\left(m\over n_{1}n_{2}\right)$
and
$\Sigma_{4}(M,N,K,\Delta,a)=\sideset{}{{}^{*}}{\sum}_{\begin{subarray}{c}n_{1},n_{2}\sim N\\ (n_{1},n_{2})=\Delta\end{subarray}}a_{n_{1}}\overline{a_{n_{2}}}\sum_{\begin{subarray}{c}m\text{ odd}\\ s(m)\leq K\end{subarray}}W\left(\frac{m}{M}\right)\left(m\over n_{1}n_{2}\right).$
2.2. Asymptotic developments
Applying Poisson summation formula to $\Sigma_{3}$ and $\Sigma_{4}$ produces asymptotic formulas:
Lemma 3.
Let $\varepsilon>0$ be arbitrary and $M,N,K,\Delta$ satisfy
$N>\Delta,\quad N^{2}M^{-1}(MN)^{\varepsilon}\leq K\leq M(MN)^{-\varepsilon}.$
Then $\Sigma_{3}=M_{3}+E_{3}\|a\|^{2}$, where
$M_{3}=\sideset{}{{}^{*}}{\sum}_{b\leq K}\sqrt{\frac{M}{b}}\sum_{\begin{subarray}{c}n_{1},n_{2}\sim N\\ (n_{1},n_{2})=\Delta\end{subarray}}a_{n_{1}}\overline{a_{n_{2}}}{\varphi(q)\over 2q}\kappa(\Delta,q)\left(\frac{b}{q}\right)\int_{0}^{+\infty}W(x^{2})\mathrm{d}x$
and
${E_{3}\over\exp_{2}(C\varepsilon^{-1})}\leq 1+(MN)^{\varepsilon}\Delta\frac{M}{N}m_{0}\sqrt{D_{1}D_{2}}\mathcal{B}\left(B,{N\over D_{1}\Delta}\right)^{\frac{1}{2}}\mathcal{B}\left(B,{N\over D_{2}\Delta}\right)^{\frac{1}{2}},$
in which
$q={n_{1}n_{2}\over\Delta^{2}},\quad\kappa(\Delta,q)=\prod_{p|\Delta}\left(1-\left(\frac{p}{q}\right)p^{-\frac{1}{2}}\right),\quad m_{0}=\min\left\{1,{N/\sqrt{MB}\over D_{1}D_{2}}\right\},$
and $B,D_{1},D_{2}$ are some numbers satisfying
(15) $D_{1},D_{2}\gg{1\over\log(2MN)},\quad 1\ll B\ll K.$
Remark.
The lower bound $K\geq N^{2}M^{-1}(MN)^{\varepsilon}$ is required to make truncations during Poisson summation.
Lemma 4.
Let $\varepsilon>0$, $N>\Delta$, and $K\leq M^{1-\varepsilon}$. Then $\Sigma_{4}=M_{4}+E_{4}\|a\|^{2}$, where
$M_{4}=\sideset{}{{}^{*}}{\sum}_{\begin{subarray}{c}v\leq K\\ (v,2\Delta)=1\end{subarray}}\sqrt{\frac{M}{v}}\sum_{\begin{subarray}{c}n_{1},n_{2}\sim N\\ (n_{1},n_{2})=\Delta\end{subarray}}a_{n_{1}}\overline{a_{n_{2}}}{\varphi(q\Delta)\over 2q\Delta}\left(\frac{v}{q}\right)\int_{0}^{+\infty}W(x^{2})\mathrm{d}x$
and
${E_{4}\over\exp_{2}(C\varepsilon^{-1})}\leq 1+(MN)^{\varepsilon}\left(M\over BD_{1}D_{2}\right)^{\frac{1}{2}}\mathcal{B}\left(B,{N\over D_{1}\Delta}\right)^{\frac{1}{2}}\mathcal{B}\left(B,{N\over D_{2}\Delta}\right)^{\frac{1}{2}},$
in which $B,D_{1},D_{2}$ are some numbers satisfying (15) and
$D_{1}D_{2}\gg\varepsilon^{2}(MN)^{-\varepsilon}\Delta^{-1}M^{\frac{1}{2}}B^{-\frac{1}{2}}.$
To produce upper bounds for $\Sigma_{2}$ (hence for $\mathcal{C}(M,N,K,\Delta)$), we estimate the difference of the main terms:
Lemma 5.
When $1\ll\Delta\ll N$ and $20<K\leq M/$ 2, there is
$M_{3}-M_{4}\leq\exp_{2}(C\varepsilon^{-1})(MN)^{\varepsilon}M^{\frac{1}{2}}K^{-\frac{1}{2}}\mathcal{B}(K\Delta^{2}L_{1},NL_{2}),$
in which $L_{1}=C_{1}\varepsilon^{-1}(MN)^{\varepsilon}$ and $L_{2}=C_{2}\varepsilon^{-1}(MN)^{\varepsilon}$ for some absolute $C_{1},C_{2}>0$.
2.3. Recursive estimates
Combining Lemma 3, Lemma 4, and Lemma 5, we find that a given estimate of $\mathcal{B}(M,N)$ can be used to produce bounds for $\mathcal{C}(M,N,K,\Delta)$. By Lemma 2, this translates into a bound for $\mathcal{B}(M,N,K)$, which is also an upper bound for $\mathcal{B}(M,N)$ according to the inequality (14), so we can estimate $\mathcal{B}(M,N)$ recursively.
If there is some $\xi>0$ and some decreasing $F(\cdot)\geq 1$ such that
(16) $\mathcal{B}(M,N)\leq F(\varepsilon)(MN)^{\varepsilon}(M+N^{\xi})$
for all $\varepsilon>0$, then our lemmas above produce a new inequality similar to (16) except with a smaller exponent $\xi$ on $N$:
Lemma 6.
Under the assumption (16), there exist absolute $C_{1},C_{2}>0$ such that
$\mathcal{C}(M,N,K,\Delta)$ $\leq\exp_{2}(C_{1}\varepsilon^{-1})F(\varepsilon)(MN)^{C_{2}\varepsilon}$
×$\times\Delta^{4}(M+N+M^{\frac{1}{2}}K^{\xi-\frac{1}{2}}+M^{\frac{1}{2}}NK^{-\frac{1}{2}}).$
Lemma 7.
Under the assumption (16), there exist absolute $C_{1},C_{2}>0$ such that
$\mathcal{B}(M,N)\leq\exp_{2}(C_{1}\varepsilon^{-1})F(\varepsilon)(MN)^{C_{2}\varepsilon^{\frac{1}{2}}}(M+M^{1-\xi}N^{2\xi-1}).$
Lemma 8.
If the assumption (16) holds for some $\xi>1$, then there exists some absolute $P>0$ such that
$\mathcal{B}(M,N)\leq\exp_{2}(P\varepsilon^{-2})F(\varepsilon^{2}P^{-1})(MN)^{\varepsilon}(M+N^{2-\frac{1}{\xi}}).$
By expanding and applying the Pólya-Vinogradov inequality, one has the trivial bound [4, p. 237]:
(17) $\mathcal{B}(M,N)\ll M+N^{2}\log N,$
which means we can launch the recursion with $\xi=2$ and $F(\varepsilon)=Q\varepsilon^{-1}$ for some large $Q>0$. Therefore, it follows from Lemma 8 that when
(18) $\xi_{r}=\begin{cases}2&r=0,\\ 2-{1\over\xi_{r-1}}&r>0\end{cases}$
and
(19) $F_{r}(\varepsilon)=\begin{cases}Q\varepsilon^{-1}&r=0,\\ \exp_{2}(P\varepsilon^{-2})F_{r-1}(\varepsilon^{2}P^{-1})&r>0,\end{cases}$
the following inequality holds for all $r\in\mathbb{N}$:
$\mathcal{B}(M,N)\leq F_{r}(\varepsilon)(MN)^{\varepsilon}(M+N^{\xi_{r}}).$
By induction, one easily finds $\xi_{r}=1+(r+1)^{-1}$, so we have
(20) $\mathcal{B}(M,N)\leq F_{\lfloor\varepsilon^{-1}\rfloor}(\varepsilon)(MN)^{2\varepsilon}(M+N).$
It remains to estimate $F_{\lfloor\varepsilon^{-1}\rfloor}(\varepsilon)$:
Proposition 1.
There exists $C_{2}>C_{1}>0$ such that for all $\varepsilon>0$ there is
$\exp_{4}(C_{1}\varepsilon^{-1})\leq F_{\lfloor\varepsilon^{-1}\rfloor}(\varepsilon)\leq\exp_{4}(C_{2}\varepsilon^{-1}).$
Plugging this into (20) and renaming $\varepsilon$ complete the proof of Theorem 1.
Remark.
The $\exp_{2}(P\varepsilon^{-2})$-factor in the recursion (19) is a consequence of estimating
$\sum_{n\sim N}[d(n)]^{k}|a_{n}|^{2}$
by $\exp_{2}(C\eta^{-1})N^{\eta}\|a\|^{2}$ via the standard bound for $d(n)$. This is optimal when $a_{n}$ is supported on primorials on which $d(n)$ is large. From this perspective, our theorem is the best one can achieve by recursive estimates with no assumptions on $\{a_{n}\}$.
3. Auxiliary lemmas
In this section, we state and prove results necessary to the treatments of $\Sigma_{3}$, $\Sigma_{4}$, and the recursive estimates. Comes first is a monotonicity principle for $\mathcal{B}(M,N)$:
Lemma 9.
There exist $C,C_{1},C_{2}>0$ such that if $M_{1},M_{2},N_{1},N_{2}$ satisfy
$M_{2}\geq M_{1}C\log(2M_{1}N_{1}),\quad N_{2}\geq N_{1}C\log(2M_{1}N_{1}),$
then
$\mathcal{B}(M_{1},N_{1})\leq C_{1}\mathcal{B}(M_{2},N_{1}),\quad\mathcal{B}(M_{1},N_{1})\leq C_{2}\mathcal{B}(M_{1},N_{2}).$
Proof.
See [4, Lemma 9]. ∎
Now, we develop a technical lemma used to estimate bilinear forms emerging in the course of estimating $E_{3}$, $E_{4}$, and $M_{3}-M_{4}$:
Lemma 10.
For complex sequences $\{a_{n}\}_{n\sim N},\{b_{n}\}_{n\sim N}$, define
(21) $S(M,N,D,a,b)=\sum_{d\sim D}\sideset{}{{}^{*}}{\sum}_{m\sim M}\Biggl{|}\sideset{}{{}^{*}}{\sum}_{\begin{subarray}{c}(n_{1},n_{2})=1\\ d|n_{1}n_{2}\end{subarray}}a_{n_{1}}b_{n_{2}}\left(m\over n_{1}n_{2}\right)\Biggr{|}$
Then we have
$S\leq\exp_{2}(C\eta^{-1})(MN)^{\eta}(D_{1}D_{2})^{\frac{1}{2}}\mathcal{B}(M,N/D_{1})^{\frac{1}{2}}\mathcal{B}(M,N/D_{2})^{\frac{1}{2}}\cdot\|a\|\cdot\|b\|,$
in which $D_{1}$ and $D_{2}$ satisfy
${1\over\log(2MN)}\ll D_{1},D_{2}\ll D,\quad D_{1}D_{2}\asymp{D\over\log^{2}(2MN)}.$
Proof.
As the result becomes trivial for large $D$, we assume $D\ll N$. Moreover, assume without loss of generality that $\{a_{n}\}$ and $\{b_{n}\}$ are supported on squarefree odd integers.
Since $(n_{1},n_{2})=1$, $d$ factors uniquely into $d_{1}d_{2}$ such that $d_{1}|n_{1}$ and $d_{2}|n_{2}$, so
$S$ $=\sum_{\begin{subarray}{c}d_{1},d_{2}\\ d_{1}d_{2}\sim D\end{subarray}}\sideset{}{{}^{*}}{\sum}_{m\sim M}\Biggl{|}\sum_{\begin{subarray}{c}(n_{1},n_{2})=1\\ d_{1}|n_{1},d_{2}|n_{2}\end{subarray}}a_{n_{1}}b_{n_{2}}\left(m\over n_{1}n_{2}\right)\Biggr{|}$
$=\sum_{\begin{subarray}{c}d_{1},d_{2}\\ d_{1}d_{2}\sim D\end{subarray}}\sideset{}{{}^{*}}{\sum}_{m\sim M}\Biggl{|}\sum_{\begin{subarray}{c}n_{1},n_{2}\\ d_{1}|n_{1},d_{2}|n_{2}\end{subarray}}a_{n_{1}}b_{n_{2}}\left(m\over n_{1}n_{2}\right)\sum_{\delta|(n_{1},n_{2})}\mu(\delta)\Biggr{|}$
$\leq\sum_{\delta\leq 2N}\sum_{\begin{subarray}{c}d_{1},d_{2}\\ d_{1}d_{2}\sim D\end{subarray}}\sideset{}{{}^{*}}{\sum}_{m\sim M}\Biggl{|}\sum_{\begin{subarray}{c}[\delta,d_{1}]|n_{1}\\ [\delta,d_{2}]|n_{2}\end{subarray}}a_{n_{1}}b_{n_{2}}\left(m\over n_{1}n_{2}\right)\Biggr{|}$
Notice that
$\sum_{\begin{subarray}{c}d_{1},d_{2}\\ d_{1}d_{2}\sim D\end{subarray}}$ $=\sum_{\mu\leq{\log D\over\log 2}}\sum_{d_{1}\sim 2^{\mu}}\sum_{d_{2}\sim D/d_{1}}\leq\sum_{\mu}\sum_{d_{1}\sim 2^{\mu}}\sum_{{D\over 2^{\mu+1}}<d_{2}\leq{D\over 2^{\mu-1}}}$
(22) $\leq\sum_{\mu}\sum_{d_{1}\sim 2^{\mu}}\sum_{\nu}\sum_{\begin{subarray}{c}{D\over 2^{\mu+1}}<d_{2}\leq{D\over 2^{\mu-1}}\\ d_{2}\sim 2^{\nu}\end{subarray}},$
so we have
$S\leq\#\{(\mu,\nu)\}\times\sum_{\delta\leq 2N}\max_{\mu,\nu}\sum_{\begin{subarray}{c}d_{1}\sim 2^{\mu}\\ d_{2}\sim 2^{\nu}\end{subarray}}\sideset{}{{}^{*}}{\sum}_{m\sim M}\Biggl{|}\sum_{\begin{subarray}{c}[\delta,d_{1}]|n_{1}\\ [\delta,d_{2}]|n_{2}\end{subarray}}a_{n_{1}}b_{n_{2}}\left(m\over n_{1}n_{2}\right)\Biggr{|}.$
By the relation (22), we see that each $\mu$ corresponds to a uniformly bounded range of $\nu$, so #$\#\{(\mu,\nu)\}=O(\log N)=O(\eta^{-1}N^{\eta/4})$, which means there exists some $P_{1},P_{2}$ such that
$1\ll P_{1},P_{2}\ll D,\quad P_{1}P_{2}\asymp D,$
and
$S\ll(\log N)\sum_{\delta\leq 2N}\sum_{\begin{subarray}{c}d_{1}\sim P_{1}\\ d_{2}\sim P_{2}\end{subarray}}\sideset{}{{}^{*}}{\sum}_{m\sim M}\Biggl{|}\sum_{\begin{subarray}{c}[\delta,d_{1}]|n_{1}\\ [\delta,d_{2}]|n_{2}\end{subarray}}a_{n_{1}}b_{n_{2}}\left(m\over n_{1}n_{2}\right)\Biggr{|}.$
Applying the Cauchy–Schwarz inequality on the right-hand side gives
(23) $S\ll(\log N)\Sigma_{a}^{\frac{1}{2}}\Sigma_{b}^{\frac{1}{2}},$
in which
$\Sigma_{a}=\sum_{\delta\leq 2N}\sum_{\begin{subarray}{c}d_{1}\sim P_{1}\\ d_{2}\sim P_{2}\end{subarray}}\sideset{}{{}^{*}}{\sum}_{m\sim M}\Biggl{|}\sum_{\begin{subarray}{c}n\\ [\delta,d_{1}]|n\end{subarray}}a_{n}\left(\frac{m}{n}\right)\Biggr{|}^{2}$
and $\Sigma_{b}$ is defined similarly. By the duality principle for bilinear forms (see [5, p. 170]) and Lemma 1, there is
$\sum_{m\sim M}\Biggl{|}\sum_{\begin{subarray}{c}n\\ [\delta,d_{1}]|n\end{subarray}}a_{n}\left(\frac{m}{n}\right)\Biggr{|}^{2}$ $\leq\mathcal{B}(N/d_{1},M)\sum_{\begin{subarray}{c}n\\ [\delta,d_{1}]|n\end{subarray}}|a_{n}|^{2}$
$\leq 2\mathcal{B}(M,N/d_{1})\sum_{\begin{subarray}{c}n\\ [\delta,d_{1}]|n\end{subarray}}|a_{n}|^{2},$
so by Lemma 9, there is some $C_{0}>0$ such that when $\theta=C_{0}\log(2MN)$, there is $\mathcal{B}(M,N/d_{1})\ll\mathcal{B}(M,\theta N/P_{1})$. Consequently, we have
$\Sigma_{a}$ $\ll P_{2}\mathcal{B}(M,\theta N/P_{1})\sum_{\delta\leq 2N}\sum_{d_{1}\sim P_{1}}\sum_{\begin{subarray}{c}n\\ [\delta,d_{1}]|n\end{subarray}}|a_{n}|^{2}$
$=P_{2}\mathcal{B}(M,\theta N/P_{1})\sum_{n\sim N}|a_{n}|^{2}\sum_{\sigma|n}\mathop{\sum_{\delta\leq 2N}\sum_{d_{1}\sim P_{1}}}_{[\delta,d_{1}]=\sigma}1.$
$\leq P_{2}\mathcal{B}(M,\theta N/P_{1})\sum_{n\sim N}|a_{n}|^{2}\sum_{\sigma|n}3^{\omega(\sigma)}$
$=P_{2}\mathcal{B}(M,\theta N/P_{1})\sum_{n\sim N}|a_{n}|^{2}4^{\omega(n)}.$
Finally, notice that when $n\sim N$ is squarefree, $4^{\omega(n)}=[d(n)]^{2}\leq\exp_{2}(C_{1}\eta^{-1})N^{\eta/2}$, so
$\Sigma_{a}\ll\exp_{2}(C_{1}\eta^{-1})N^{\eta/2}P_{2}\mathcal{B}(M,\theta N/P_{1})\|a\|^{2}.$
By the same line of reasoning, we obtain a similar bound for $\Sigma_{b}$. Plugging these bounds into (23) and setting $D_{1}=P_{1}/\theta,D_{2}=P_{2}/\theta$ complete the proof. ∎
The next few lemmas are dedicated to Poisson summations:
Lemma 11.
Let $W$ be a smooth function whose support is a compact subset of $(0,+\infty)$. Then for any positive squarefree odd $q$, there is
$\sum_{m\geq 1}W\left(\frac{m}{M}\right)\left(\frac{m}{q}\right)={M\tau(q)\over q}\sum_{h\in\mathbb{Z}}\widehat{W}\left(h\over q/M\right)\left(\frac{h}{q}\right),$
where $\tau(q)$ denotes the quadratic Gauss sum with modulus $q$.
Proof.
See [4, Lemma 11]. ∎
The following lemma is used to separate variables during the treatment of $E_{3}$:
Lemma 12.
Let $\varrho:\mathbb{R}\to\mathbb{R}$ be a Schwartz function. Define $\varrho_{+}(s),\varrho_{-}(s)$ by
$\varrho_{\pm}(s)=\int_{0}^{+\infty}\varrho(\pm x)x^{s-1}\mathrm{d}x.$
Then $\varrho_{+}(s),\varrho_{-}(s)$ are holomorphic in $\Re(s)>0$ and satisfy
$\varrho_{+}(s),\varrho_{-}(s)\ll_{A,\sigma}|t|^{-A}\quad(s=\sigma+it).$
In addition, for any $\sigma>0$, we have
$\varrho(\pm x)={1\over 2\pi i}\int_{\sigma-i\infty}^{\sigma+i\infty}x^{-s}\varrho_{\pm}(s)\mathrm{d}s,$
and when $\sigma\to 0^{+}$, we have
$\int_{-\infty}^{\infty}|\varrho_{\pm}(\sigma+it)|\mathrm{d}t\ll\sigma^{-1}.$
Proof.
Every result apart from the last one is identical to those in [4, Lemma 12]. For the last result, notice that integration by parts gives $\rho_{\pm}(s)\ll(\sigma+|t|^{2})^{-1}$. ∎
The following variant of the Poisson summation formula helps identify the main terms and error terms in $\Sigma_{3}$ and $\Sigma_{4}$:
Lemma 13.
Let $\psi:\mathbb{R}\to\mathbb{R}$ be a Schwartz function and $G,H:\mathbb{[}0,+\infty)\to\mathbb{[}0,+\infty)$ be increasing functions such that
$|\psi(x)|\leq G(A)|x|^{-A},\quad|\widehat{\psi}(x)|\leq H(A)|x|^{-A}.$
If $A\geq 0$, $0<X_{1}\leq X\leq X_{2}$, $L\geq(X_{2}/X)^{2}$ and $J=\min(X/X_{1},X/X_{2})$, then for any integer $k\geq 2$, there is
(24) $\sum_{\begin{subarray}{c}n\in\mathbb{Z}\\ (n,k)=1\end{subarray}}\psi\left(\frac{n}{X}\right)$ $={\varphi(k)\over k}X\int_{-\infty}^{+\infty}\psi(x)\mathrm{d}x-\psi(0)\sum_{\begin{subarray}{c}d|k\\ d\leq X_{2}\end{subarray}}\mu(d)$
$-\sum_{\begin{subarray}{c}d|k\\ d>X_{2}\end{subarray}}\mu(d)\frac{X}{d}\int_{-\infty}^{+\infty}\psi(x)\mathrm{d}x$
$+\sum_{1\leq|l|\leq L}\sum_{\begin{subarray}{c}d|k\\ X_{1}<d\leq X_{2}\end{subarray}}\mu(d)\frac{X}{d}\widehat{\psi}\left(lX\over d\right)$
$+\max\{G(A+1),H(A+2)\}\cdot O(XJ^{-A}).$
Proof.
By tracking the implied constants in the proof of [4, Lemma 13], one finds that the error term is of the form
$\ll G(A_{1})X(X_{2}/X)^{1-A_{1}}+H(A_{2})X(X/X_{1})^{-A_{2}}+H(A_{3})X(X_{2}/X)^{2-A_{3}},$
so setting $A_{1}=A+1$, $A_{2}=A$, and $A_{3}=A+2$ gives the desired result. ∎
The next few results concern explicit estimates for smooth weight functions. Comes first is a chain rule for higher-order derivatives:
Proposition 2 (Faà di Bruno).
Let $f(x)$ be and $g(x)$ be $n$-times differentiable. Then
${\mathrm{d}^{n}\over\mathrm{d}x^{n}}[f(g(x))]=\sum_{\begin{subarray}{c}r_{1},r_{2},\dots,r_{n}\geq 0\\ r_{1}+2r_{2}+\cdots+nr_{n}=n\end{subarray}}{n!\over r_{1}!r_{2}!\cdots r_{n}!}f^{(n)}(g(x))\prod_{j=1}^{n}\left[g^{(j)}(x)\over j!\right]^{r_{j}}.$
In particular, when $g(x)$ is a quadratic polynomial, there is
${\mathrm{d}^{n}\over\mathrm{d}x^{n}}[f(g(x))]=\sum_{0\leq r\leq n/2}{n!\over 2^{r}r!(n-2r)!}f^{(n-r)}(g(x))[g^{\prime}(x)]^{n-2r}[g^{\prime\prime}(x)]^{r}.$
Proof.
For the general case, see [10, Theorem 2]. The following is a proof for the case where $g(x)$ is a quadratic polynomial:
When $n=0$, there is nothing to prove. Suppose the result holds at $n$. Since
${\mathrm{d}\over\mathrm{d}x}$$\left\{f^{(n-r)}(g(x))[g^{\prime}(x)]^{n-2r}[g^{\prime\prime}(x)]^{r}\right\}$
$=f^{(n+1-r)}(g(x))[g^{\prime}(x)]^{n+1-2r}[g^{\prime\prime}(x)]^{r}$
$+(n-2r)f^{(n-r)}(g(x))[g^{\prime}(x)]^{n-1-2r}[g^{\prime\prime}(x)]^{r+1},$
writing $n=2\mu+\nu$ for $\nu\in\{0,1\}$ gives
(25) ${\mathrm{d}^{n+1}\over\mathrm{d}x^{n+1}}[f(g(x))]$ $=f^{(n+1)}(g(x))[g^{\prime}(x)]^{n+1}$
(26) $+\sum_{r=1}^{\mu}{n!\over 2^{r}r!(n-2r)!}f^{(n+1-r)}(g(x))[g^{\prime}(x)]^{n+1-2r}[g^{\prime\prime}(x)]^{r}$
(27) $+\sum_{r=0}^{\mu-1}{n!\over 2^{r}r!(n-1-2r)!}f^{(n-r)}(g(x))[g^{\prime}(x)]^{n-1-2r}[g^{\prime\prime}(x)]^{r+1}$
(28) $+{n!(n-2\mu)\over 2^{\mu}\mu!(n-2\mu)!}f^{(n-\mu)}(g(x))[g^{\prime}(x)]^{n-1-2\mu}[g^{\prime\prime}(x)]^{\mu+1}.$
By re-indexing the term (27), we compute that the sum of (26) and (27) is
$\sum_{r=1}^{\mu}\left[{n!\over 2^{r}r!(n-2r)!}+{n!\over 2^{r-1}(r-1)!(n+1-2r)!}\right]$
×$\times f^{(n+1-r)}(g(x))[g^{\prime}(x)]^{n+1-2r}[g^{\prime\prime}(x)]^{r}$
$=\sum_{r=1}^{\mu}{(n+1)!\over 2^{r}r!(n+1-2r)!}f^{(n+1-r)}(g(x))[g^{\prime}(x)]^{n+1-2r}[g^{\prime\prime}(x)]^{r}.$
Extending this sum to $r=0$ includes the right-hand side of (25). When $\nu=0$, (28) vanishes. When $\nu=1$, (28) becomes
$={n!\over 2^{\mu}\mu!}f^{(n-\mu)}(g(x))[g^{\prime}(x)]^{n-1-2\mu}[g^{\prime\prime}(x)]^{\mu+1}$
$={n!(\mu+1)\over 2^{\mu}(\mu+1)!}f^{(n+1-(\mu+1))}(g(x))[g^{\prime}(x)]^{n+1-2(\mu+1)}[g^{\prime\prime}(x)]^{\mu+1}.$
Notice that $\mu+1=(n+1)/2=\lfloor(n+1)/2\rfloor$, so in all cases, the combination of (25)-(28) is
$=\sum_{0\leq r\leq(n+1)/2}{(n+1)!\over 2^{r}r!(n+1-2r)!}f^{(n+1-r)}(g(x))[g^{\prime}(x)]^{n+1-2r}[g^{\prime\prime}(x)]^{r},$
completing the proof. ∎
Proposition 3.
Let $W(x)$ be as in (13). Then there exist absolute constants $C_{1},C_{2},C_{3}>0$ satisfying
$|W^{(j)}(x)|\leq C_{1}^{j}j^{3j},\quad|\widehat{W}^{(j)}(k)|\leq C_{2}^{j}C_{3}^{A}A^{3A}|k|^{-A}$
for all $A,j\geq 1$.
Proof.
Let $f(y)=e^{y}$ and $g(x)={-1\over 1-x}$. Define $h(x)=f(g(x))$ and
$\omega(x)=e^{-{2\over 1-x^{2}}}\chi_{[-1,1]}(x)=h(x)\cdot h(-x)\chi_{[-1,1]}(x),$
so there is $W(x)=e^{\frac{8}{3}}\omega\left(x-\frac{3}{2}\right)$. Since for $10\leq x<$ 1,
$g^{(j)}(x)=(-1)(-2)\cdots(-j)\cdot{1\over(x-1)^{j+1}}=-j![-g(x)]^{j+1}.$
By Proposition 2, we deduce that for $10\leq x<$ 1,
$|h^{(n)}(x)|$ $\leq\sum_{\begin{subarray}{c}r_{1},r_{2},\dots,r_{n}\geq 0\\ r_{1}+2r_{2}+\dots+nr_{n}=n\end{subarray}}{n!\over r_{1}!r_{2}!\cdots r_{n}!}e^{g(x)}\prod_{j=1}^{n}[-g(x)]^{(j+1)r_{j}}$
$\leq\sum_{\begin{subarray}{c}r_{1},r_{2},\dots,r_{n}\geq 0\\ r_{1}+2r_{2}+\dots+nr_{n}=n\end{subarray}}{n!\over r_{1}!r_{2}!\cdots r_{n}!}e^{g(x)}\prod_{j=1}^{n}[-g(x)]^{2jr_{j}}$
$=[-g(x)]^{2n}e^{g(x)}\sum_{\begin{subarray}{c}r_{1},r_{2},\dots,r_{n}\geq 0\\ r_{1}+2r_{2}+\dots+nr_{n}=n\end{subarray}}{n!\over r_{1}!r_{2}!\cdots r_{n}!}\leq[-g(x)]^{2n}e^{g(x)}n!e^{n}.$
Finally, by the elementary inequality $u^{R}e^{-u}\leq R^{R}e^{-R}$ for all $R,u>0$, we have
$|h^{(n)}(x)|\leq(2n)^{2n}n!e^{-n}\leq 4^{n}n^{3n}.$
By the Leibniz rule for derivatives, this means
$|\omega^{(j)}(x)|$ $\leq\sum_{n=0}^{j}\binom{j}{n}[4^{n}n^{3n}]\cdot[4^{j-n}(j-n)^{3(j-n)}]$
$\leq 4^{j}\sum_{n=0}^{j}\binom{j}{n}[j^{3n}]\cdot[j^{3(j-n)}]\leq 8^{j}j^{3j}.$
Thus, we can take $C_{1}=8e^{\frac{8}{3}}$ in the estimate for $W^{(j)}(x)$. The estimate for $\widehat{W}^{(j)}(k)$ is a consequence of integration by parts. ∎
The following propositions are used to establish explicit asymptotic estimates for the approximate functional equation for $L(s,\chi)^{2}$:
Proposition 4.
Let $s=\sigma+it$ with $\sigma$ bounded and $|t|\geq 1$ and $u=x+iy$ with $x\to+\infty$. Then
${\Gamma(s+u)\over\Gamma(s)}\ll C^{x}x^{2x}|t|^{x}e^{\frac{\pi}{2}|y|}.$
uniformly for all $y\in\mathbb{R}$ and $|t|\geq 1$.
Proof.
By Stirling’s approximation, when $|\eta|\ll|\xi|$, there is
$|\Gamma(\eta+i\xi)|=|\xi|^{\eta-\frac{1}{2}}e^{-\frac{\pi}{2}|\xi|}e^{O(|\eta|)}.$
Combining with the fact that $|\Gamma(\eta+i\xi)|\leq\Gamma(\eta)\ll\eta^{\eta}$ for $\eta>0$, we have
${\Gamma(s+u)\over\Gamma(s)}$ $\ll C_{1}^{x}x^{x}{(|t+y|+1)^{\sigma+x-\frac{1}{2}}e^{-\frac{\pi}{2}|t+y|}\over|t|^{\sigma-\frac{1}{2}}e^{-\frac{\pi}{2}|t|}}$
$=C_{1}^{x}x^{x}|t|^{\frac{1}{2}-\sigma}(|t+y|+1)^{x+\sigma-\frac{1}{2}}e^{\frac{\pi}{2}(|t|-|t+y|)}.$
If $|y|\leq 4|t|$, it follows from $|t|-|t+y|\leq|y|$ that
$(|t+y|+1)^{x+\sigma-\frac{1}{2}}e^{\frac{\pi}{2}(|t|-|t+y|)}\ll C_{2}^{x}|t|^{x+\sigma-\frac{1}{2}}e^{\frac{\pi}{2}|y|}.$
If $|y|>4|t|$, then it follows from $|t|-|t+y|\leq 2|t|-|y|<-|y|/2$ that
$(|t+y|+1)$ ${}^{x+\sigma-\frac{1}{2}}e^{\frac{\pi}{2}(|t|-|t+y|)}\ll C_{2}^{x}|t|^{x+\sigma-\frac{1}{2}}e^{\frac{\pi}{2}|y|}$
$\ll C_{3}^{x}(\pi|y|/4)^{x+\sigma-\frac{1}{2}}e^{-\pi|y|/4}.$
$\leq C_{3}^{x}\left(x+\sigma-\frac{1}{2}\right)^{x+\sigma-\frac{1}{2}}\ll C_{4}^{x}x^{x}.$
Combining the bounds from both cases gives the desired result. ∎
Proposition 5.
If $0\leq\sigma\leq 1$, $|t|\geq 1$, and $a\in\{0,1\}$. Define $s=\sigma+it$ and
(29) $V_{s}(y)={1\over 2\pi i}\int_{3-i\infty}^{3+i\infty}(\pi y)^{-u}e^{u^{2}}\left\{\Gamma\left(s+u+a\over 2\right)\over\Gamma\left(s+a\over 2\right)\right\}^{2}{\mathrm{d}u\over u}.$
Then for all $A\geq 1$, there is
(30) $V_{s}(y)\ll C^{A}A^{A}e^{A^{2}}(y/|t|)^{-A}$
as $y\to+\infty$ and
(31) $V_{s}(y)=1+O\{(y/|t|)^{1/3}\}$
as $y\to 0^{+}$. The implied constants are absolute.
Remark.
This is an explicit form of [5, Proposition 5.4] particularized for $L(s,\chi)^{2}$.
Proof.
To prove the bound (30), we move the line of integration to $\Re(u)=A$, so writing $u=A+iv$, Proposition 4 gives
$\left\{\Gamma\left(s+u+a\over 2\right)\over\Gamma\left(s+a\over 2\right)\right\}^{2}\ll C_{1}^{A}A^{A}|t|^{A}e^{\frac{\pi}{2}|y|}.$
Notice that $|e^{u^{2}}|=e^{A^{2}-v^{2}}$, so (29) becomes
$V_{s}(y)\ll C_{1}^{A}A^{A}(\pi y)^{-A}|t|^{A}e^{A^{2}}\int_{-\infty}^{\infty}{e^{\frac{\pi}{2}|v|-v^{2}}\over|v|+1}\mathrm{d}v\ll C_{2}^{A}A^{A}e^{A^{2}}(y/|t|)^{-A}.$
For the asymptotic (31), observe that the integrand has a simple pole at $u=0$ with residue $1$, so moving the line of integration to $\Re(u)=-\frac{1}{3}$ and proceed similarly give us the desired result. ∎
4. Proof of Theorem 1
This section proves the results stated in Section 2. For the sake of bookkeeping, Section 4.X gives demonstrations of propositions in Section 2.X.
4.1. Preparation for Poisson summation
Proof of Lemma 2.
Most of Heath-Brown’s arguments can be reused. The $\varepsilon$-factor is a consequence of plugging the divisor bound $d(n)\leq\exp_{2}(C\varepsilon^{-1})n^{\varepsilon}$ into
$\mathcal{B}^{*}(M,K)\sum_{n\sim N}d(n)^{2}|a_{n}|^{2}$
of [4, p. 254]. ∎
4.2. Asymptotic developments
In this section, we prove Lemma 3, Lemma 4, and Lemma 5.
Proof of Lemma 3.
According to [4, p. 254], applying Lemma 11 to $\Sigma_{3}$ gives
$\Sigma_{3}=\sum_{e|2\Delta}{M\mu(e)\over e}\sum_{a}\sum_{c\geq 1}\sideset{}{{}^{*}}{\sum}_{\begin{subarray}{c}n_{1},n_{2}\sim N/\Delta\\ (n_{1},n_{2})=1\end{subarray}}a_{n_{1}}^{(1)}\overline{a}_{n_{2}}^{(1)}\sideset{}{{}^{*}}{\sum}_{b\geq 1}\widehat{W}\left(abc^{2}\over en_{1}n_{2}/M\right)\left(b\over n_{1}n_{2}\right),$
in which $\sum_{a}$ sums over ±$\pm 1,\pm 2$ and
$a_{n}^{(1)}=a_{n\Delta}{\tau(n)\over n}\left(eac^{2}\over n\right).$
Then Heath-Brown shows that when $N^{2}M^{-1}(MN)^{\varepsilon}\leq K\leq M(MN)^{-\varepsilon}$, the portion of $\Sigma_{3}$ with $b>K$ is $\ll_{\varepsilon}\|a\|^{2}$. Now, we give an explicit estimate for the $\ll_{\varepsilon}$ constant.
According to Proposition 3, we have $\widehat{W}(x)\ll\exp(A^{2})|x|^{-A}$, so the terms with $b>K$ satisfies
$\ll\exp(A^{2})K(\Delta KMN^{-2})^{-A}\sum_{e|2\Delta}{M\over e}\sum_{a}\sum_{c\geq 1}\sideset{}{{}^{*}}{\sum}_{\begin{subarray}{c}n_{1},n_{2}\sim N/\Delta\\ (n_{1},n_{2})=1\end{subarray}}|a_{n_{1}}^{(1)}\overline{a}_{n_{2}}^{(1)}|c^{-2A}$
$\ll\exp(A^{2})KM(\Delta KMN^{-2})^{-A}\sum_{e\leq 2\Delta}{1\over e}\sideset{}{{}^{*}}{\sum}_{\begin{subarray}{c}n_{1},n_{2}\sim N/\Delta\end{subarray}}|a_{n_{1}\Delta}a_{n_{2}\Delta}|$
$\ll\exp(A^{2})M^{2}(MN)^{-A\varepsilon}{\log 2\Delta\over\Delta^{A}}\sum_{n_{1},n_{2}\sim N}|a_{n_{1}}a_{n_{2}}|.$
Since $|\alpha\beta|\leq(|\alpha|^{2}+|\beta|^{2})/2$, we have
$\sum_{n_{1},n_{2}\sim N}|a_{n_{1}}a_{n_{2}}|\leq\frac{1}{2}\sum_{n_{1},n_{2}\sim N}(|a_{n_{1}}|^{2}+|a_{n_{2}}|^{2})=N\|a\|^{2},$
so setting $A=2\varepsilon^{-1}$ means the contribution of terms with $b>K$ is majorized by
(32) $\ll\exp(A^{2})M^{2}N(MN)^{-A\varepsilon}\|a\|^{2}\ll\exp(4\varepsilon^{-2})\|a\|^{2},$
which is sufficient for our purpose as $\exp(4\varepsilon^{-2})\ll\exp_{2}(\varepsilon^{-1})$.
For terms in $\Sigma_{3}$ satisfying $b\leq K$, Heath-Brown first divides the range $b\leq K$ into $b\sim B$ (totally $O(\log M)$ subintervals). The terms of $\Sigma_{3}$ within $b\sim B$ can be written as
$\Sigma_{3}^{(B)}=\sum_{e|2\Delta}{M\mu(e)\over e}\sum_{a}\sideset{}{{}^{*}}{\sum}_{b\sim B}\sideset{}{{}^{*}}{\sum}_{\begin{subarray}{c}n_{1},n_{2}\sim N/\Delta\\ (n_{1},n_{2})=1\end{subarray}}a_{n_{1}}^{(2)}\overline{a}_{n_{2}}^{(2)}S(n_{1}n_{2}),$
where
$a_{n}^{(2)}=a_{n\Delta}{\tau(n)\over n}\left(ea\over n\right)$
and
$S(q)=\frac{1}{2}\sum_{\begin{subarray}{c}c\in\mathbb{Z}\\ (c,q)=1\end{subarray}}\widehat{W}\left(abc^{2}\over eq/M\right).$
We are now in a position to apply Lemma 13 with
(33) $X=\sqrt{eq\over Mb},\quad\psi(x)=\widehat{W}(ax^{2}),$
$X_{1}=(MN)^{-\eta/2}{\sqrt{e}\over\Delta}N(MB)^{-\frac{1}{2}},\quad X_{2}=(MN)^{\eta}{\sqrt{e}\over\Delta}N(MB)^{-\frac{1}{2}},$
in which $\eta>0$ is arbitrary, allowing us to take $J=(MN)^{\eta/4}$ and $L=(MN)^{3\eta}$, so $S(q)$ becomes
$S(q)=S_{1}(q)-S_{2}(q)-S_{3}(q)+S_{4}(q)+R(q),$
where the $S_{k}(q)$ corresponds to the $k$’th main term in the expansion (24) and $R(q)$ corresponds to the error term.
Using the exact arguments in [4, p. 255], collecting all the $S_{1}$-portion in $\Sigma_{3}^{(B)}$ produces the main term $M_{3}$. Terms arising from $S_{2}(q)$, $S_{3}(q)$, and $S_{4}(q)$ can be bounded using Heath-Brown’s original reasoning except that each application of [4, Lemma 10] and [4, Lemma 12] is replaced with our explicit Lemma 10 and Lemma 12. These terms eventually produce the estimate of $E_{3}$ in the statement of Lemma 3.
For the $R$-portion of $\Sigma_{3}^{(B)}$, we need $G(A)$ and $H(A)$ for our choice of $\psi(x)$ in (33). By Proposition 3, we can take $G(A)=C_{1}^{A}A^{3A}$ for some large fixed $C_{1}>0$. As for $H(A)$, plugging $f(x)=\widehat{W}(x)$ and $g(x)=ax^{2}$ into Proposition 2, so
$|\psi^{(n)}(x)|$ $\leq\sum_{0\leq r\leq n/2}{n!\over 2^{r}r!(n-2r)!}|\widehat{W}^{(n-r)}(ax^{2})|\cdot|2ax|^{n-2r}|2a|^{r}$
$\leq\sum_{0\leq r\leq n/2}{n!\over 2^{r}r!(n-2r)!}|\widehat{W}^{(n-r)}(ax^{2})|\cdot(4|x|)^{n}$
$\leq(4|x|)^{n}\cdot\sup_{n/2\leq j\leq n}|\widehat{W}^{(j)}(ax^{2})|\cdot\sum_{0\leq r\leq n/2}{n!\over 2^{r}r!}$
$\leq e^{\frac{1}{2}}n!(4|x|)^{n}\cdot\sup_{n/2\leq j\leq n}|\widehat{W}^{(j)}(ax^{2})|.$
Now, using the bounds for $\widehat{W}^{(j)}$ in Proposition 3 and setting $A=n+2$, we have
$|\psi^{(n)}(x)|\leq n^{2n}C_{3}^{A}A^{3A}|x|^{-(A-n)}\leq C_{4}^{n}n^{9n}|x|^{-2}.$
Through integration by parts, we see that $H(A)=C_{5}^{A}A^{9A}$ is admissible. Similar to how we deduce the bound (32), we set $A=C_{6}\varepsilon^{-1}$ in the error term of (24), so the error term of $\Sigma_{3}$ arised from $R(q)$ is majorized by $\exp_{2}(\varepsilon^{-1})\|a\|^{2}$. ∎
Proof of Lemma 4.
Similar to the proof of Lemma 3, we apply Proposition 3 and Proposition 2 so that when $\psi(x)=W(x^{2})$, we can take $G(A),H(A)=C_{1}^{A}A^{C_{2}A}$ for some absolute $C_{1},C_{2}>0$. Therefore, replacing the application of [4, Lemma 10] and [4, Lemma 13] in Heath-Brown’s derivation with our explicit Lemma 10 and Lemma 13 gives us the desired $M_{4}$ and $E_{4}$. ∎
Proof of Lemma 5.
Most of Heath-Brown’s original arguments can be re-used except we apply the explicit divisor bound $d(w)\ll\exp_{2}(C\eta^{-1})w^{\eta}$ and replace the use of [4, Lemma 10] with our Lemma 10. ∎
4.3. Recursive estimates
In this section, we prove Lemma 6, Lemma 7, and Lemma 8, which are needed to perform recursive estimates. We also derive Proposition 1 which analyzes the growth $F_{r}(\varepsilon)$ mentioned in Section 2.3.
Proof of Lemma 6.
The $F(\varepsilon)$-factor comes from [4, (20)]:
$\mathcal{B}\left(B,{N\over D_{1}\Delta}\right)^{\frac{1}{2}}\mathcal{B}\left(B,{N\over D_{2}\Delta}\right)^{\frac{1}{2}}$
$\ll F(\varepsilon)\left(B\cdot{N\over D_{1}}\right)^{\varepsilon/2}\left(B\cdot{N\over D_{2}}\right)^{\varepsilon/2}(B^{\xi}+N/D_{1})^{\frac{1}{2}}(B^{\xi}+N/D_{2})^{\frac{1}{2}}$
$\ll F(\varepsilon)(BN)^{\varepsilon}(B^{\xi}+N^{\frac{1}{2}}B^{\xi/2}+ND_{1}^{-\frac{1}{2}}D_{2}^{-\frac{1}{2}})$
$\ll F(\varepsilon)(MN)^{\varepsilon}(B^{\xi}+N^{\frac{1}{2}}B^{\xi/2}+ND_{1}^{-\frac{1}{2}}D_{2}^{-\frac{1}{2}}),$
and $\mathcal{B}(K\Delta^{2}L_{1},NL_{2})$ term in Lemma 5 with $L_{1}=C_{1}\varepsilon^{-1}(MN)^{\varepsilon}$ and $L_{2}=C_{2}\varepsilon^{-1}(MN)^{\varepsilon}$ (see also the fifth equation in [4, p. 264]):
$\mathcal{B}(K\Delta^{2}L_{1},NL_{2})$ $\leq F(\varepsilon)(K\Delta^{2}L_{1}NL_{2})^{\varepsilon}\{(K\Delta^{2}L_{1})^{\xi}+NL_{2}\}$
$\leq\exp_{2}(C_{3}\varepsilon^{-1})F(\varepsilon)(MN)^{C_{4}\varepsilon}\Delta^{4}(M^{\frac{1}{2}}K^{\xi-\frac{1}{2}}+M^{\frac{1}{2}}NK^{-\frac{1}{2}}),$
where the $\Delta^{4}$ factor follows from the fact that $\xi\leq 2$. Plugging these two explicit estimates into Heath-Brown’s arguments produces the desired explicit upper bound for $\mathcal{C}(M,N,K,\Delta)$ in Lemma 6. ∎
Proof of Lemma 7.
Replacing [4, Lemma 2] and [4, Lemma 6] with our explicit Lemma 2 and Lemma 6, the arguments in [4, p. 264] lead to an explicit form of [4, (24)]). That is, whenever $1\leq\Delta_{0}<N\leq N_{0}$,
(34) $N_{0}(MN_{0})^{\varepsilon}\leq M,$
and $K=N_{0}^{2}M^{-1}(MN_{0})^{\varepsilon}$, there is
$\mathcal{B}(M,N,K)\leq G\{\mathcal{B}(M,N_{1},K)+H\},$
in which $N_{1}\leq N\Delta_{0}^{-1}$,
$G=\exp_{2}(C_{1}\varepsilon^{-1})(MN_{0})^{C_{2}\varepsilon}\geq 2,$
and
$H=F(\varepsilon)\Delta_{0}^{5}(M+M^{1-\xi}N_{0}^{2\xi-1}).$
By the nature of Lemma 2, there is a sequence $N_{r}$ satisfying $N_{r}\leq N_{r-1}\Delta_{0}^{-1}$ and
$\mathcal{B}(M,N_{r-1},K)\leq G\{\mathcal{B}(M,N_{r},K)+H\}.$
In other words, we have
$\mathcal{B}(M,N_{r-1},K)+{GH\over G-1}\leq G\left\{\mathcal{B}(M,N_{r},K)+{GH\over G-1}\right\},$
so there is
$\mathcal{B}(M,N_{0},K)\leq G^{r}\{\mathcal{B}(M,N_{r},K)+GH\}\leq G^{2r}\{\mathcal{B}(M,N_{r},K)+H\}.$
Let $\Delta_{0}=N_{0}^{1/R}$, so there is some $1\leq r_{0}\leq R$ such that $N_{r_{0}}\leq\Delta_{0}<N_{r_{0}-1}$, allowing us to conclude that
$\mathcal{B}(M,N_{0},K)$ $\leq\exp_{2}(C_{1}\varepsilon^{-1}+\log 2r_{0})(MN_{0})^{2C_{2}r_{0}\varepsilon}\{\mathcal{B}(M,N_{r_{0}},K)$
$+N_{0}^{5/R}F(\varepsilon)(M+M^{1-\xi}N_{0}^{2\xi-1})\}.$
Plugging in the trivial bound $\mathcal{B}(M,N_{r_{0}},K)\ll MN_{r_{0}}\leq MN_{0}^{1/R}$ gives
$\mathcal{B}(M,N_{0},K)\leq\exp_{2}(C_{1}\varepsilon^{-1}+\log 2R)F(\varepsilon)(MN_{0})^{2C_{2}R\varepsilon+\frac{5}{R}}(M+M^{1-\xi}N_{0}^{2\xi-1}).$
Let $R=\varepsilon^{-\frac{1}{2}}$, $C_{3}>2C_{1}$, and $C_{4}>2C_{2}$, so we have
$\mathcal{B}(M,N_{0},K)\leq\exp_{2}(C_{3}\varepsilon^{-1})(MN_{0})^{C_{4}\varepsilon^{\frac{1}{2}}}F(\varepsilon)(M+M^{1-\xi}N_{0}^{2\xi-1}),$
Since $\Sigma_{1}\leq\Sigma_{1}^{(K)}$, this is an upper bound for $\mathcal{B}(M,N_{0})$ as well, completing the proof in the case under the assumption (34). For the case where this assumption is false, it follows directly from the estimate (16) and $\xi\leq 2$ that
$\mathcal{B}(M,N_{0})$ $\leq F(\varepsilon)(MN_{0})^{\varepsilon}(M+N_{0}^{\xi})$
$=F(\varepsilon)(MN_{0})^{\varepsilon}(M+M^{1-\xi}N_{0}^{2\xi-1}(M/N_{0})^{\xi-1})$
$\leq F(\varepsilon)(MN_{0})^{2\varepsilon}(M+M^{1-\xi}N_{0}^{2\xi-1}),$
which is consistent with the conclusion as $2\varepsilon\leq C_{4}\varepsilon^{\frac{1}{2}}$. ∎
Proof of Lemma 8.
Let $\alpha>0$ be some parameter to be determined later. If $M\leq N^{\alpha}$, then it follows from Lemma 9 and Lemma 7 that there is some $C_{0}>0$ satisfying
$\mathcal{B}(M,N)$ $\ll\mathcal{B}(N^{\alpha}C_{0}\varepsilon^{-\frac{1}{2}}(MN)^{\varepsilon^{\frac{1}{2}}},N)$
$\ll\exp_{2}(C_{1}\varepsilon^{-1})F(\varepsilon)[\varepsilon^{-\frac{1}{2}}N^{\alpha+1}(MN)^{\varepsilon^{\frac{1}{2}}}]^{C_{2}\varepsilon^{\frac{1}{2}}}(N^{\alpha}+N^{\alpha(1-\xi)+2\xi-1})$
$\leq\exp_{2}(C_{3}\varepsilon^{-1})F(\varepsilon)(MN)^{C_{4}\varepsilon^{\frac{1}{2}}}(N^{\alpha}+N^{\alpha(1-\xi)+2\xi-1}).$
Let $\alpha=2-\xi^{-1}$ so that both exponents in the last line are equal. Suppose $M\geq N^{\alpha}$. Then it follows from Lemma 7 that
$\mathcal{B}(M,N)$ $\leq\exp_{2}(C_{1}\varepsilon^{-1})F(\varepsilon)(MN)^{C_{2}\varepsilon^{\frac{1}{2}}}(M+N^{\alpha(1-\xi)+2\xi-1})$
$=\exp_{2}(C_{1}\varepsilon^{-1})F(\varepsilon)(MN)^{C_{2}\varepsilon^{\frac{1}{2}}}(M+N^{2-\xi^{-1}}).$
Therefore, in any case, there is
$\mathcal{B}(M,N)\leq\exp_{2}(C_{3}\varepsilon^{-1})F(\varepsilon)(MN)^{C_{4}\varepsilon^{\frac{1}{2}}}(M+N^{2-\xi^{-1}}),$
provided that $C_{3}>C_{1}$. Let $Q=\max(C_{4}^{2},1)$, so there is
$\mathcal{B}(M,N)\leq\exp_{2}(C_{3}Q^{2}\varepsilon^{-2})F(\varepsilon^{2}/Q)(MN)^{\varepsilon}(M+N^{2-\xi^{-1}}).$
Setting $P=C_{3}Q^{2}>Q$ and using the monotonicity of $F(\cdot)$ produce the conclusion. ∎
Proof of Proposition 1.
Define $\varepsilon_{0}=\varepsilon$ and $\varepsilon_{r}=\varepsilon_{r-1}^{2}/P$. Then $\varepsilon_{r}=P^{1-2^{r}}\varepsilon^{2^{r}}$ and it follows from the definition (19) of $F_{r}(\varepsilon)$ that
(35) $F_{r}(\varepsilon)=Q\exp(e^{1/\varepsilon_{1}}+e^{1/\varepsilon_{2}}+\dots+e^{1/\varepsilon_{r}})\varepsilon_{r}^{-1},$
which means when $r=\lfloor\varepsilon\rfloor$, there is
$F_{r}(\varepsilon)$ $\leq Q\exp(re^{1/\varepsilon_{r}})\varepsilon_{r}^{-1}\leq Q\exp\left(\varepsilon^{-1}e^{P^{2^{\varepsilon^{-1}}-1}\varepsilon^{-2^{\varepsilon^{-1}}}}\right)P^{2^{\varepsilon^{-1}}-1}\varepsilon^{-2^{\varepsilon^{-1}}}$
$=Q\exp_{2}\left(P^{2^{\varepsilon^{-1}}-1}e^{2^{\varepsilon^{-1}}\log\varepsilon^{-1}}+\log\varepsilon^{-1}\right)P^{2^{\varepsilon^{-1}}-1}e^{2^{\varepsilon^{-1}}\log\varepsilon^{-1}}$
$\leq Q\exp_{2}(\exp_{2}(C_{1}\varepsilon^{-1})+\log\varepsilon^{-1})\exp_{2}(C_{2}\varepsilon^{-1})\leq\exp_{4}(C_{3}\varepsilon^{-1}).$
On the other side, the expression (35) can be bounded below using
$F_{r}(\varepsilon)$ $\geq\exp_{2}(1/\varepsilon_{r})\geq\exp_{3}\{(2^{r}-1)(\log P+\log\varepsilon^{-1})\}$
$\geq\exp_{3}(2^{r}C_{4})\geq\exp_{3}(2^{\varepsilon^{-1}-1}C_{4})\geq\exp_{4}(C_{5}\varepsilon^{-1}),$
completing the proof. ∎
5. Applications
In this section, we apply Theorem 1 to derive the explicit quadratic large sieve inequality (Corollary 1). Subsequently, we apply this result to moments of Dirichlet $L$-functions (Corollary 2), and rational primes splitting in quadratic fields (Corollary 3).
5.1. Explicit quadratic large sieve inequality
To prove Corollary 1 from Theorem 1, we begin by proving an explicit form of the inequality (10) by breaking the range of $m$ and $n$ into dyadic intervals so that for $M,N\geq 2$
$\sideset{}{{}^{*}}{\sum}_{m\leq M}\left|\sideset{}{{}^{*}}{\sum}_{n\leq N}a_{n}\left(\frac{n}{m}\right)\right|^{2}$ $\ll(\log M)\max_{1\leq M^{\prime}\leq M}\sum_{m\sim M^{\prime}}\left|\sideset{}{{}^{*}}{\sum}_{n\leq N}a_{n}\left(\frac{n}{m}\right)\right|^{2}$
$\ll(\log M)(\log N)^{2}\max_{\begin{subarray}{c}1\leq M^{\prime}\leq M\\ 1\leq N^{\prime}\leq N\end{subarray}}\sum_{m\sim M^{\prime}}\left|\sideset{}{{}^{*}}{\sum}_{n\sim N^{\prime}}a_{n}\left(\frac{n}{m}\right)\right|^{2}$
$\leq(\log MN)^{3}\max_{\begin{subarray}{c}1\leq M^{\prime}\leq M\\ 1\leq N^{\prime}\leq N\end{subarray}}\mathcal{B}(M^{\prime},N^{\prime})\sum_{n\sim N^{\prime}}|a_{n}|^{2}.$
Notice that $\log MN\ll\varepsilon^{-1}(MN)^{\varepsilon/2}$, so plugging in Theorem 1 with $\varepsilon$ replaced with $\varepsilon/2$ gives
(36) $\sideset{}{{}^{*}}{\sum}_{m\leq M}\left|\sideset{}{{}^{*}}{\sum}_{n\leq N}a_{n}\left(\frac{n}{m}\right)\right|^{2}\leq\exp_{4}(C\varepsilon^{-1})(MN)^{\varepsilon}(M+N)\|a\|^{2}.$
For the inequality (2), we can reuse the arguments in [4, p. 265–267]. To obtain explicit $\varepsilon$-dependence in $\ll$-constants, it suffices to notice that all $X^{\varepsilon}$-terms during the derivation are introduced only through the use of (36), the divisor bound $d(n)\leq\exp_{2}(C\varepsilon^{-1})n^{\varepsilon}$, and $\log X\ll\varepsilon^{-1}X^{\varepsilon}$. Thus, we have
(37) $\sum_{\chi\in S(Q)}\left|\sum_{n\leq N}a_{n}\chi(n)\right|^{2}\leq\exp_{4}(C\varepsilon^{-1})(QN)^{\varepsilon}(Q+N)\sum_{\begin{subarray}{c}n_{1},n_{2}\leq N\\ n_{1}n_{2}=\square\end{subarray}}|a_{n_{1}}a_{n_{2}}|,$
which becomes Corollary 1 after setting $\varepsilon=C^{\prime}/\log_{4}(QN)$ for some large $C^{\prime}>0$.
5.2. Moments of Dirichlet $L$-functions
In this section, we prove Corollary 2 by showing that when $s=\frac{1}{2}+it$, $T=|t|+1$, and $S^{\prime}(Q)=S(2Q)\smallsetminus S(Q)$ represents all primitive quadratic characters $\chi$ modulo $q_{\chi}\sim Q$, there is
(38) $S(Q,s)=\sum_{\chi\in S^{\prime}(Q)}|L(s,\chi)|^{4}\ll\exp_{4}(C\varepsilon^{-1})(QT)^{1+\varepsilon},$
which immediately implies (4) by dyadic summation.
Heath-Brown’s original argument began with the formula
$L(s,\chi)^{2}=\sum_{n\geq 1}{d(n)\chi(n)\over n^{s}}e^{-n/U}-{1\over 2\pi i}\int_{\alpha-i\infty}^{\alpha+i\infty}U^{w-s}\Gamma(w-s)L(w,\chi)^{2}\mathrm{d}w$
for some $0<\alpha<\frac{1}{2}$ and used a recursive estimate process to obtain the bound
(39) $S(Q,\sigma+it)\ll_{\sigma,\varepsilon}(QT)^{\varepsilon}(Q+(QT)^{2-2\sigma})$
valid for all fixed $\sigma>\frac{1}{2}$. Explicating Heath-Brown’s argument indeed produces (38). However, as demonstrated in Section 4.3, tracking $\ll$-constants in recursive estimates is a complicated procedure. In this section, we give a simpler proof of (38) by appealing to the approximate functional equation, which completely avoids the need for recursion.
For primitive quadratic $\chi$ modulo $q=q_{\chi}$, it follows from [5, §5.2] that
$L(s,\chi)^{2}=\sum_{n\geq 1}{d(n)\chi(n)\over n^{\frac{1}{2}+it}}V_{\frac{1}{2}+it}\left(n\over q\right)+\varepsilon(t)\sum_{n\geq 1}{d(n)\chi(n)\over n^{\frac{1}{2}-it}}V_{\frac{1}{2}-it}\left(n\over q\right),$
where $V_{s}(y)$ is defined in (29) and $\varepsilon(t)$ is some function satisfying $|\varepsilon(t)|=1$.
Applying Cauchy–Schwartz inequality, we obtain
$S(Q,s)\leq 2S_{1}(Q,t)+2S_{1}(Q,-t),$
where
$S_{1}(Q,t)=\sum_{\chi\in S^{\prime}(Q)}\left|\sum_{n\geq 1}{d(n)\chi(n)\over n^{\frac{1}{2}+it}}V_{\frac{1}{2}+it}\left(n\over q_{\chi}\right)\right|^{2}.$
By Proposition 5, when $X\gg QT$ and $A\gg 1$, we have
$\sum_{n>X}{d(n)\chi(n)\over n^{\frac{1}{2}+it}}V_{\frac{1}{2}+it}\left(n\over q_{\chi}\right)\ll e^{2A^{2}}X(QTX^{-1})^{A}.$
Setting $0<\delta\leq 1$, $X=(QT)^{1+\delta}$, and $A=2\delta^{-1}$ makes this term $\ll\exp(8\delta^{-2})$, so one has
$S_{1}(Q,t)$ $\ll Q\exp(16\delta^{-2})+\sum_{\chi\in S^{\prime}(Q)}\left|\sum_{n\leq X}{d(n)\chi(n)\over n^{\frac{1}{2}+it}}V_{\frac{1}{2}+it}\left(n\over q_{\chi}\right)\right|^{2}.$
$\ll Q\exp(16\delta^{-2})+(\log X)^{2}\max_{1\leq N\leq X}\sum_{\chi\in S^{\prime}(Q)}\left|\sum_{n\sim N}{d(n)\chi(n)\over n^{\frac{1}{2}+it}}V_{\frac{1}{2}+it}\left(n\over q_{\chi}\right)\right|^{2}.$
By (37) and the fact that $V_{s}(y)=O(1)$, we deduce that
$\sum_{\chi\in S^{\prime}(Q)}$ $\left|\sum_{n\sim N}{d(n)\chi(n)\over n^{\frac{1}{2}+it}}V_{\frac{1}{2}+it}\left(n\over q_{\chi}\right)\right|^{2}$
$\ll\exp_{4}(C_{1}\varepsilon^{-1})(QN)^{\varepsilon/8}(Q+N)N\max_{n\sim N}|d(n)^{2}n^{-1}|$
$\ll\exp_{4}(C_{2}\varepsilon^{-1})(QN)^{\varepsilon/4}(Q+N)\leq\exp_{4}(C_{2}\varepsilon^{-1})(QX)^{\varepsilon/4}(Q+X)$
$\ll\exp_{4}(C_{2}\varepsilon^{-1})(QT)^{1+{3\varepsilon\over 4}+\delta}.$
Finally, set $\delta=\varepsilon/8$, so it follows from $\log X\ll\varepsilon^{-1}(QT)^{\varepsilon/16}$ and $\exp(16\delta^{-2})\ll\exp_{4}(C_{2}\varepsilon^{-1})$ that
$S_{1}(Q,t)\ll\exp_{4}(C_{3}\varepsilon^{-1})(QT)^{1+\varepsilon}.$
Plugging this into $S(Q,s)$ gives the desired result.
5.3. Rational primes splitting in $\mathbb{Q}(\sqrt{-q})$
In this section, we prove Corollary 3. Let $q$ denote a prime $\equiv 3\pmod{4}$.
By a well known property of quadratic fields [9, Proposition 8.5], a rational prime $p$ splits in $\mathbb{Q}(\sqrt{-q})$ if and only if
$\left(-q\over p\right)=1,$
so it follows from the reciprocity law that the number of such $p\leq X$ is
$P_{q}(X)=\frac{1}{2}\sum_{\begin{subarray}{c}2<p\leq X\\ p\neq q\end{subarray}}\left[1+\left(\frac{p}{q}\right)\right].$
Now, let $m_{Q}$ be the number of $q\sim Q$ such that $|P_{q}(X)-\frac{1}{2}\pi(X)|>\delta\pi(X)$. Then there is
$m_{Q}\delta^{2}\pi(X)^{2}\leq\frac{1}{4}\sum_{q\sim Q}\left|\sum_{\begin{subarray}{c}2<p\leq X\\ p\neq q\end{subarray}}\left(\frac{p}{q}\right)\right|^{2},$
so it follows from Theorem 1 that
$m_{Q}\ll\delta^{-2}\exp_{4}\left\{C\log(XQ)\over\log_{4}(XQ)\right\}\left(1+\frac{Q}{X}\right).$
Finally, we set $\log X\geq 3C\log Q/\log_{4}Q$ so that $\log X\geq 2C\log(XQ)/\log_{4}(XQ)$, completing the proof of Corollary 3.